To get an idea of the time required for diffusion, estimate how long it might take for ammonia (NH3) to be detected 10 cm from a bottle after it is opened, assuming only diffusion is occurring.
Answer #1

time required for diffusion is calculated by

t = x^2/2D

x = mean distance traveled by the diffusing solute

D = diffusion coefficient

for Ammonia D = 1.51*10^-5 cm^2/sec

using given data

t = (10 cm)^2/(2*1.51*10^-5 cm^2/sec)

t = 3.311*10^6 sec = 38.32 days